# Score identity Distillation: Exponentially Fast Distillation of Pretrained Diffusion Models for One-Step Generation

### v1

@@ 4.3 Loss Approximation via Projected Score Matching
4.4 Fused Loss of SiD
${\mathcal{L}}_{\theta}^{(2)}=\penalty 10000\$${\mathcal{L}}_{\theta}^{(1)}+\mathbb{E}_{\bm{x}_{g}\sim p_{\theta}(\bm{x}_{g})}\mathbb{E}_{\bm{x}_{t}\sim q(\bm{x}_{t}\,|\,\bm{x}_{g},t)}$
$\left[{\sigma_{t}^{-2}}{{\delta}_{\phi,\psi}(\bm{x}_{t})^{T}(f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g})}\right].$ (21)
$\hat{{\mathcal{L}}}_{\theta}^{\triangle}=\hat{{\mathcal{L}}}_{\theta}^{(2)}-\hat{{\mathcal{L}}}_{\theta}^{(1)}={\sigma_{t}^{-2}}{\delta}_{\phi,\psi}(\bm{x}_{t})^{T}(f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g})$
directs the gradient towards the desired direction.
Thus we are propelled to consider the loss
In summary, the weighted loss is expressed as
$\textstyle\tilde{L}_{\theta}(\bm{x}_{t},t,\phi,\psi)=(1-\alpha)\frac{\omega(t)}{{\sigma_{t}^{4}}}\|f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t)\|_{2}^{2}$
$\textstyle+\frac{\omega(t)}{{\sigma_{t}^{4}}}(f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t))^{T}(f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g}),$ (23)
where $\bm{x}_{t}$ is generated as in (18) and $\omega(t)$ are weighted coefficients that need to be specified. To compute the gradient of the above equation, SiD backpropagates the gradient through both $\phi$ and $\psi$ by calculating two score gradients ($i.e.$, gradients of scores) as
$\textstyle\nabla_{\theta}f_{\phi}(\bm{x}_{t},t)=\frac{\partial f_{\phi}(\bm{x}_{t},t)}{\partial\bm{x}_{t}}\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\bm{z})$ (24)
$\textstyle\nabla_{\theta}f_{\psi}(\bm{x}_{t},t)=\frac{\partial f_{\psi}(\bm{x}_{t},t)}{\partial\bm{x}_{t}}\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\bm{z}).$
This feature distinguishes SiD from Diff-Instruct and DMD that do not use score gradients $\frac{\partial f_{\phi}(\mathbf{x}_{t},t)}{\partial\bm{x}_{t}}$ and $\frac{\partial f_{\psi}(\bm{x}_{t},t)}{\partial\bm{x}_{t}}$.
4.5 Noise Weighting and Scheduling
$\omega(t)=C\sigma_{t}^{4}/\|\bm{x}_{g}-f_{\phi}(\bm{x}_{t},t)\|_{1,sg}\penalty 10000\ \penalty 10000\ .$ (25)
The distillation process is outlined in Algorithm˜1. The one-step generation procedure is straightforward:
$\bm{x}=G_{\theta}(\sigma_{\text{init}}\bm{z}),\penalty 10000\ \bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
where $\sigma_{\text{init}}$, by default set to 2.5, remains consistent throughout distillation and generation.
@@ 4.5 Noise Weighting and Scheduling
   Sample $\bm{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\bm{x}_{g}=G_{\theta}(\sigma_{\text{init}}\bm{z})$
   Sample $t\sim p(t),\penalty 10000\ \bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$ and $\bm{x}_{t}=\bm{x}_{g}+\sigma_{t}\bm{\epsilon}_{t}$
   Update $\psi$ with Equation˜10:
    $\mathcal{L}_{\psi}={\gamma(t)}\|f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g}\|_{2}^{2}$
@@
   Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$ with Equation˜26, compute $\omega_{t}$ with Equation˜25, and update $G_{\theta}$ with Equation˜23:
    $\tilde{\mathcal{L}}_{\theta}=(1-\alpha)\frac{\omega(t)}{{\sigma_{t}^{4}}}\|f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t)\|_{2}^{2}$
    $\frac{\omega(t)}{{\sigma_{t}^{4}}}(f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t))^{T}(f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g})$
@@
Table 1: Comparison of various deep generative models trained on CIFAR-10 without label conditioning. The best and second-best one/few-step generators under the FID or IS metric are highlighted with bold and italic bold, respectively.
[…]
Figure 4: Evolution of FIDs for the SiD generator during the distillation of the EDM teacher model pretrained on CIFAR-10 (unconditional), using $\alpha=1.0$ or $\alpha=1.2$ and a batch size of 256. The performance of EDM, along with DMD and Diff-Instruct, is depicted with horizontal lines in purple, green, and red, respectively.
5 Experimental Results
Table 2: Analogous to Table 1 for CIFAR-10 (conditional).
[…]
Figure 5: Analogous to Fig. 4 for CIFAR-10 (conditional).
Implementation details. We implement SiD based on the EDM (Karras et al., 2022) code base and we initialize both the generator $G_{\theta}$ and its score estimation network $f_{\psi}$ by copying the architecture and parameters of the pretrained score network $f_{\phi}$ from EDM (Karras et al., 2022). We provide the other implementation details in Appendix˜C.
Table 3: Analogous to Table 1 for ImageNet 64x64 with label conditioning. The Precision and Recall metrics are also included.
[…]
Figure 6: Analogous plot to Fig. 4 for ImageNet 64x64. The batch size is 8192. See the results of batch size 1024 in Fig. 9.
Table 4: Analogous to Table 1 for FFHQ 64x64.
Figure 7: Analogous plot to Fig. 4 for FFHQ 64x64. The batch size is 512.
Table 5: Analogous to Table 1 for AFHQ-v2 64x64.
[…]
Figure 8: Ananagous plot to Fig. 4 for AFHQ-v2 64x64. The batch size is 512.
5.1 Ablation Study and Parameter Settings
The results indicate a stable performance of the model when $\alpha$ varies from 0 to 1.2. A negative value of $\alpha$ results in large FIDs. This observation supports our analysis in section˜4.2 that directly optimizing $\mathcal{L}_{\theta}^{(1)}$ given by (13) may not lead to meaningful improvement, as our loss, shown in (22), is $\mathcal{L}_{\theta}^{(2)}-\alpha\mathcal{L}_{\theta}^{(1)}$. As $\alpha$ increases within the tested range, we observe a gradual improvement in IS and FID performance, peaking at $\alpha=1$ or $\alpha=1.2$. Based on these findings, we select $\alpha=1$ or $\alpha=1.2$ for all our experiments, although a more refined grid search on $\alpha$ might reveal even better performance outcomes.
Instead, we increased the overall batch size from 1024 to 8192 (while keeping the batch per GPU unchanged at 16, requiring more gradient accumulation rounds) and reduced the learning rate from 5e-6 to 4e-6. Under $\alpha=1$, we observed stable performance, while under $\alpha=1.2$, we observed occasional spikes in FID. Upon examining the generations corresponding to these spikes, as shown in the fifth image of Fig. 11 and Fig. 12 in the Appendix, we found interesting patterns where certain uncommon features, such as nests containing birds, were exaggerated. However, with a batch size as large as 8192, these occasional spikes did not seem to significantly impact the overall declining trend, which was roughly log-log linear initially and gradually leveled off. With that said, when the batch size was reduced to 1024, the sudden divergence could potentially be caused by such a spike, as observed in Fig. 9.
The drawback of using a larger batch size in this case is that it takes SiD longer to outperform Diff-instruct and DMD, as clearly shown by comparing the FID trajectories in Figs. 6 and 9. Although it’s feasible to develop more advanced strategies, including progressively increasing the batch size, annealing the learning rate, and implementing gradient clipping, we’ll reserve these for future study.
5.2 Benchmark Performance
Our comprehensive evaluation compares SiD against leading deep generative models, encompassing both distilled diffusion models and those built from scratch. Random images generated by SiD in a single step are displayed in Figs. 13-16 in the Appendix.
The comparative analysis, detailed in Tables 1-5 and illustrated in Figs. 4-8, underlines the single-step SiD generator’s proficiency in leveraging the insights from the pretrained EDM (teacher diffusion model) across a variety of benchmarks, including CIFAR-10 (both conditional and unconditional formats), ImageNet 64x64, FFHQ 64x64, and AFHQ-v2 64x64. Remarkably, the SiD-trained generator surpasses the EDM teacher in nearly all tested environments, showcasing its enhanced performance not just relative to the original multi-step teacher model but also against a broad spectrum of cutting-edge models, from traditional multi-step diffusion models to the latest single-step distilled models and GANs. The sole deviation in this pattern occurs with ImageNet 64x64, where SiD, at $\alpha=1.2$, attains an FID of 1.524, which is exceeded by Jabri et al. (2022)’s RIN at 1.23 FID with 1000 steps and the teacher model VP-EDM’s 1.36 FID with 511 steps.
Convergence Speed. Our SiD generator, designed for distilling pretrained diffusion models, rapidly achieves the capability to generate photo-realistic images in a single step. This efficiency is showcased in Fig. 1 for the EDM model pretrained on ImageNet 64x64 and in Fig. 2 for CIFAR 32x32 (unconditional). The performance of the SiD method is further highlighted in Figs. 4-8, where the x-axis represents the thousands of images processed during training. These figures track the FID’s evolution across four datasets for both $\alpha=1$ and $\alpha=1.2$, demonstrating a roughly linear relationship between the logarithm of the FID and the logarithm of the number of processed images. This relationship indicates that FID decreases exponentially as distillation progresses, a trend that is observed or expected to eventually slow down and approach a steady state.
6 Conclusion
Impact Statement
Acknowledgments
The authors would like to express their gratitude to Dr. Zhuoran Yang for his valuable comments and suggestions.
Appendix A Proofs
@@
For Gaussian diffusion, we have (2), which we explore to derive the identity shown below. While $p_{\theta}(\bm{x}_{t})$ often does not have an analytic form, exploiting its semi-implicit construction, its score can be expressed as
$\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})$ $=\frac{\int\nabla_{\bm{x}_{t}}q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{{\theta}}(\bm{x}_{g})d\bm{x}_{g}}{p_{\theta}(\bm{x}_{t})}$
$=\frac{\int q(\bm{x}_{t}\,|\,\bm{x}_{g})\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})d\bm{x}_{g}}{p_{\theta}(\bm{x}_{t})}$
$=-\frac{\int q(\bm{x}_{t}\,|\,\bm{x}_{g})\frac{\bm{x}_{t}-a_{t}\bm{x}_{g}}{\sigma_{t}^{2}}p_{\theta}(\bm{x}_{g})d\bm{x}_{g}}{p_{\theta}(\bm{x}_{t})}$
$=-\frac{\bm{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma_{t}^{2}}\frac{\int\bm{x}_{g}q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})d\bm{x}_{g}}{p_{\theta}(\bm{x}_{t})}$
$=-\frac{\bm{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma^{2}}\int\bm{x}_{g}q(\bm{x}_{g}\,|\,\bm{x}_{t})d\bm{x}_{g}$
$=-\frac{\bm{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma_{t}^{2}}\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]$ (27)
$\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})$ $=\frac{\int\nabla_{\bm{x}_{t}}q(\bm{x}_{t}\,|\,\bm{x}_{0})p_{{\phi}}(\bm{x}_{0})d\bm{x}_{0}}{p_{\phi}(\bm{x}_{t})}$ (28)
@@
$\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]=\frac{\bm{x}_{t}+\sigma_{t}^{2}\nabla_{\bm{x}_{t}}\ln q_{g}(\bm{x}_{t})}{a_{t}}$ (29)
@@
$\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[u^{T}(\bm{x}_{t})\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})]$
$=\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[u^{T}(\bm{x}_{t})\frac{\nabla_{\bm{x}_{t}}p_{\theta}(\bm{x}_{t})}{p_{\theta}(\bm{x}_{t})}]$
$=\int{u^{T}(\bm{x}_{t})\nabla_{\bm{x}_{t}}p_{\theta}(\bm{x}_{t})}d\bm{x}_{t}$
$=\int u^{T}(\bm{x}_{t})\int\nabla_{\bm{x}_{t}}q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})d\bm{x}_{g}d\bm{x}_{t}$
$=\int u^{T}(\bm{x}_{t})\int q(\bm{x}_{t}\,|\,\bm{x}_{g})\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})d\bm{x}_{g}d\bm{x}_{t}$
$=\mathbb{E}_{(\bm{x}_{t},\bm{x}_{g})\sim q(\bm{x}_{t}\,|\,\bm{x}_{g})p_{\theta}(\bm{x}_{g})}[u^{T}(\bm{x}_{t})\nabla_{\bm{x}_{t}}\ln q(\bm{x}_{t}\,|\,\bm{x}_{g})].$ (30)
@@
$\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}\|S(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})\|_{2}^{2}$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[(\mathbb{E}[\bm{x}_{0}\,|\,\bm{x}_{t}]-\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}])^{T}(S(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t}))]$
@@
① $=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}(\mathbb{E}[\bm{x}_{0}\,|\,\bm{x}_{t}]-\bm{x}_{t})]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\mathbb{E}[\bm{x}_{0}\,|\,\bm{x}_{t}]]-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\bm{x}_{t})]$ (32)
② $=\mathbb{E}_{\bm{x}_{g}\sim p_{\theta}(\bm{x}_{g})}\mathbb{E}_{\bm{x}_{t}\sim q(\bm{x}_{t}\,|\,\bm{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t}\,|\,\bm{x}_{g}))]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{g}\sim p_{\theta}(\bm{x}_{g})}\mathbb{E}_{\bm{x}_{t}\sim q(\bm{x}_{t}\,|\,\bm{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}(\bm{x}_{g}-\bm{x}_{t})]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{g}\sim p_{\theta}(\bm{x}_{g})}\mathbb{E}_{\bm{x}_{t}\sim q(\bm{x}_{t}\,|\,\bm{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\bm{x}_{g}]-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\bm{x}_{t}]$ (33)
@@
Appendix B Analytic study of the toy example
We prove the conclusions in Propositions˜4 and 6. Given $p_{\text{data}}(\bm{x}_{0})=\mathcal{N}(\mathbf{0},\mathbf{I})$, $p_{\text{data}}(\bm{x}_{t})=\mathcal{N}(0,1+\sigma_{t}^{2})$, and $q(\bm{x}_{t}\,|\,\bm{x}_{g})=\mathcal{N}(\bm{x}_{g},\sigma_{t}^{2}),\penalty 10000\ p_{\theta}(\bm{x}_{g})=\mathcal{N}(\theta,1)$, the optimal value of $\theta$ would be $\theta^{*}=0$. The score
$S(\bm{x}_{t})=\nabla_{\bm{x}_{t}}\ln p_{\text{data}}(\bm{x}_{t})=-\frac{\bm{x}_{t}}{1+\sigma_{t}^{2}}$
$\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})=-\frac{\bm{x}_{t}-\theta}{1+\sigma_{t}^{2}}.$
Hence, the difference between the scores is $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})=-\frac{\theta}{1+\sigma_{t}^{2}}$. By Tweedie’s formula
$f_{\phi}(\bm{x}_{t},t)=\mathbb{E}[\bm{x}_{0}\,|\,\bm{x}_{t}]=\frac{\bm{x}_{t}}{1+\sigma_{t}^{2}}$
$\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]=\bm{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\theta\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}$

### v2

@@ Appendix A Ablation Study and Parameter Settings
Appendix B Algorithm Box
@@
   Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$
   Sample $t\sim p(t),~\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$ and $\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
   Update $\psi$ with Equation 10:
       $\mathcal{L}_{\psi}={\gamma(t)}\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}$
@@
   Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$ with Equation 26, compute $\omega_{t}$ with Equation 25, and update $G_{\theta}$ with Equation 23:
       $\tilde{\mathcal{L}}_{\theta}=(1-\alpha)\frac{\omega(t)}{{\sigma_{t}^{4}}}\|f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t)\|_{2}^{2}$
             $\frac{\omega(t)}{{\sigma_{t}^{4}}}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t))^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})$
@@
Appendix C Proofs
@@
For Gaussian diffusion, we have (2), which we explore to derive the identity shown below. While $p_{\theta}(\boldsymbol{x}_{t})$ often does not have an analytic form, exploiting its semi-implicit construction, its score can be expressed as
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$ $=\frac{\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{{\theta}}(\boldsymbol{x}_{g})d\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
$=\frac{\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
$=-\frac{\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\frac{\boldsymbol{x}_{t}-a_{t}\boldsymbol{x}_{g}}{\sigma_{t}^{2}}p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
$=-\frac{\boldsymbol{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma_{t}^{2}}\frac{\int\boldsymbol{x}_{g}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
$=-\frac{\boldsymbol{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma^{2}}\int\boldsymbol{x}_{g}q(\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t})\mathrm{d}\boldsymbol{x}_{g}$
$=-\frac{\boldsymbol{x}_{t}}{\sigma_{t}^{2}}+\frac{a_{t}}{\sigma_{t}^{2}}\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]$ (27)
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})$ $=\frac{\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})p_{{\phi}}(\boldsymbol{x}_{0})\mathrm{d}\boldsymbol{x}_{0}}{p_{\phi}(\boldsymbol{x}_{t})}$ (28)
@@
$\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]=\frac{\boldsymbol{x}_{t}+\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\ln q_{g}(\boldsymbol{x}_{t})}{a_{t}}$ (29)
@@
$\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[u^{T}(\boldsymbol{x}_{t})\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})]$
$=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[u^{T}(\boldsymbol{x}_{t})\frac{\nabla_{\boldsymbol{x}_{t}}p_{\theta}(\boldsymbol{x}_{t})}{p_{\theta}(\boldsymbol{x}_{t})}]$
$=\int{u^{T}(\boldsymbol{x}_{t})\nabla_{\boldsymbol{x}_{t}}p_{\theta}(\boldsymbol{x}_{t})}\mathrm{d}\boldsymbol{x}_{t}$
$=\int u^{T}(\boldsymbol{x}_{t})\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{\mathrm{}}d\boldsymbol{x}_{g}d\boldsymbol{x}_{t}$
$=\int u^{T}(\boldsymbol{x}_{t})\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}d\boldsymbol{x}_{t}$
$=\mathbb{E}_{(\boldsymbol{x}_{t},\boldsymbol{x}_{g})\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})}[u^{T}(\boldsymbol{x}_{t})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})].$ (30)
@@
$\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\|S(\boldsymbol{x}_{t})-\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})\|_{2}^{2}$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[(\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]-\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}])^{T}(S(\boldsymbol{x}_{t})-\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t}))]$
@@
① $=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]-\boldsymbol{x}_{t})]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]]-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\boldsymbol{x}_{t})]$ (32)
② $=\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g}))]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(\boldsymbol{x}_{g}-\boldsymbol{x}_{t})]$
$=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\boldsymbol{x}_{g}]-\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\boldsymbol{x}_{t}]$ (33)
@@
Appendix D Analytic study of the toy example
We prove the conclusions in Propositions 4 and 6. Given $p_{\text{data}}(\boldsymbol{x}_{0})=\mathcal{N}(\mathbf{0},\mathbf{I})$, $p_{\text{data}}(\boldsymbol{x}_{t})=\mathcal{N}(0,1+\sigma_{t}^{2})$, and $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})=\mathcal{N}(\boldsymbol{x}_{g},\sigma_{t}^{2}),~p_{\theta}(\boldsymbol{x}_{g})=\mathcal{N}(\theta,1)$, the optimal value of $\theta$ would be $\theta^{*}=0$. The score
$S(\boldsymbol{x}_{t})=\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})=-\frac{\boldsymbol{x}_{t}}{1+\sigma_{t}^{2}}$
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})=-\frac{\boldsymbol{x}_{t}-\theta}{1+\sigma_{t}^{2}}.$
Hence, the difference between the scores is $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})=-\frac{\theta}{1+\sigma_{t}^{2}}$. By Tweedie’s formula
$f_{\phi}(\boldsymbol{x}_{t},t)=\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]=\frac{\boldsymbol{x}_{t}}{1+\sigma_{t}^{2}}$
$\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]=\boldsymbol{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\theta\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}$